# Revising Multimodal VAEs with Diffusion Decoders

### v1

@@ 3.3 Conditional Sampling
To perform conditional sampling (i.e. generate missing modalities given some modalities) utilizing our model, we can leverage the factorized nature of the posterior distribution. We first sample from the joint posterior distribution conditioned on the available modalities and then use the sampled latent variable to generate samples using the respective generative model for each modality. The posterior will be the product of the observed modalities $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})=\prod_{i\in{o}}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where the subscript $o$ denotes observed modalities. Then, we can generate an unobserved modality ${u}$ from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})$ by feeding it to the appropriate decoder.
@@ 3.4 Unconditional Sampling
Unconditional generation happens when no specific modality is provided as input. In this scenario, the goal is to produce samples that exhibit coherence across all modalities without relying on any initial conditions. Previous multimodal VAEs perform unconditional generation by first sampling from a $\mathcal{N}(0,\mathbf{I})$ and then feeding the sample to each decoder. Doing that may produce suboptimal results as there is a gap between the posterior and the prior (Aneja et al., 2021). To get better samples, we train an auxiliary score-based diffusion model to sample from the posterior. Figure  2 demonstrates how the score model facilitates unconditional sampling. This auxiliary score model is trained using the diffusion objective but on the latent space $\mathbf{z}$ to generate samples from the posterior. The objective is $\mathbb{E}_{t,\mathbf{z}_{0},\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t)\right\|^{2}$ where $\mathbf{z_{0}}$ is sampled from $q_{\phi}(\mathbf{z}|\mathbf{x})=\prod_{i}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where $\mathbf{x}_{i}$ consists of each modality in $\mathbf{x}$. The main objective of training this auxiliary model is to get high-quality samples from the posterior instead of using $\mathcal{N}(0,\mathbf{I})$ as the starting sample for unconditional generation. To generate samples unconditionally, we first sample $\mathbf{z}$ from the auxiliary score model and use the respective generative model $p_{\theta}(\mathbf{x}_{i}|\mathbf{z})$ of each modality to get the final samples.
@@ 4 Experiments
We perform experiments on two datasets. The first one is from Wah et al. (2011), which contains images of birds with text captions describing the birds. We use actual images of the birds instead of ResNET features, unlike previous multimodal VAEs that struggle with image data (Shi et al., 2019). The second dataset is a high-dimensional CelebAMask-HQ  (Lee et al., 2020) dataset that consists of images, masks, and attributes of celebrities expressed in the three modalities. We compare different multimodal VAE baselines consisting of MVTCAE by Hwang et al. (2021), MoPoE by Sutter et al. (2021), and MMVAE+ Palumbo et al. (2023). MMVAE+ is not included in the CUB dataset experiment as we use pre-trained VAE for the text captions, which will require changing the structure of the text VAE because of the modality-specific and shared representations in MMVAE+. Nonetheless, we added the results from their paper for the text-to-image results, which can be referenced from their work (Palumbo et al., 2023). We also add additional baselines from  Wesego and Rooshenas (2023); Bounoua et al. (2024) that use a latent score-based model to learn the joint latent space discussed in the next sections.
@@ 4.1 Caltech Birds (CUB)
The CUB dataset consists of two modalities: image and text describing the image (caption). In order to get meaningful text outputs, we initialized the text VAE from the pre-trained weight of  Li et al. (2020) that uses a BERT encoder and GPT-2 decoder for all models. The images are resized to 64x64. We used a latent size of $768$ for both the image and text. The encoder architecture for the image modality is similar to that of  Daniel and Tamar (2021) and the diffusion decoder uses a UNET architecture. The models were trained for 500 epochs using the Adam optimizer (Kingma and Ba, 2015). The baseline models are trained using different $\beta$ values listed in the appendix, and the best model was selected by using the average of the conditional and unconditional FID. The unconditional auxiliary score model also uses a 1D UNET architecture and accepts input similar to the size of the latent dimension. DDIM sampling was used for 50 steps to generate samples for the image modality and for unconditional sampling (Song et al., 2021). More details about the training and inference are in the Appendix section.
@@
this bird has a pointed bill with a yellow breast and belly
Figure 4: Conditional generation given a text using MDDVAE
@@
We use FID, which is a widely adopted metric for evaluating the quality of images, as the evaluation metric to measure conditionally and unconditionally generated samples (Heusel et al., 2017). In addition to that, we use Clip-Score to measure the similarity of the image-text pairs even if it’s not the best metric for this dataset  (Hessel et al., 2021; Bounoua et al., 2024). It’s clear that a good multimodal VAE model will have coherent unconditional samples where the image and text represent the same concept expressed in its respective modality. In addition, conditionally generated images should follow the description in the text, and conditionally generated text should follow the given image.
bird has gray body feathers white breast feather and black beak
a small with a white belly brown breast and crown and a yellow throat
Figure 6: Unconditional generation using MDDVAE
We present the results in Table  1 and  2. In addition to the implemented baselines, we add results from previous works of MLD by  Bounoua et al. (2024) that uses latent diffusion model and MMVAE+ on text-to-image FID. The FID scores evaluation demonstrates that our model, MDDVAE, outperforms the baselines, generating high-quality images that exhibit greater coherence and alignment with the text modality. In contrast, the baselines struggle to generate meaningful images that are consistent with the textual information, supporting the findings of Daunhawer et al. (2022) regarding the challenges faced by existing multimodal VAEs in difficult tasks. Table  2 also shows that MDDVAE generates more coherent conditional outputs that follow the given modality. We also added qualitative results from our model, which is the only multimodal VAE model generating good images in this dataset, are shown in figure  4 and  6 for conditional and unconditional generation respectively.
@@ 4.2 CelebAMask-HQ
The CelebAMask-HQ dataset provides multimodal representations of visual characteristics related to human faces. In this dataset, the images, masks, and attributes can be viewed as distinct yet complementary modalities that capture various aspects of an individual’s appearance. We resize the images and masks to a resolution of 128 by 128 pixels. We combine multiple masks of parts of the face into a single 0/1 mask similar to  Wesego and Rooshenas (2023). Specifically, all the provided masks in the CelebAMask-HQ dataset, except for the skin mask, are drawn on top of each other, resulting in a single composite mask. This composite mask encodes the presence or absence of various facial features. For the attribute modality, we follow the preprocessing approach described by Wu and Goodman (2018); Wesego and Rooshenas (2023). Out of the 40 existing attributes in the CelebAMask-HQ dataset, we selectively choose 18 attributes that are most relevant and informative for representing the visual characteristics of a person’s face. We used a latent size of $256$. Similar to the previous dataset, the encoder and decoder architecture for the image modality are inspired by  Daniel and Tamar (2021) except for the diffusion decoder, which uses a UNET architecture. The unconditional auxiliary score-based model also works on the $256$ dimension latent size resized to 16x16 utilizing a UNET architecture. Sampling is done by using the DDIM sampling technique for 50 steps (Song et al., 2021). More details about the experimental setup are located in the appendix section. The baselines here are MoPoE, MVTCAE, MMVAE+ plus SBM-VAE-C and SBM-RAE-C by  Wesego and Rooshenas (2023) and MLD by  Bounoua et al. (2024), which use unimodal variational autoencoders (VAEs) or autoencoders (AEs) for each modality and fuse them using a score-based model for learning the joint distribution.
(a) MDDVAE
(b) MVTCAE
(c) MMVAE+
(d) MoPoE
@@
MDDVAE 0.76($\pm$0.001) 0.76($\pm$0.001) 32.4($\pm$0.08) 32.4($\pm$0.06) 28.3($\pm$0.27) 32.4($\pm$0.25) 0.90($\pm$0.001) 0.90($\pm$0.001)
@@
We present our results in table  3 which shows the performance of the models in both conditional and unconditional settings. MDDVAE generates high-quality images, achieving the best FID scores among all models in image generation. Not only that, but having a better representation model in the image modality also helps the performance of the model in the other modalities where MDDVAE outperforms the baselines in attribute prediction, even if the attribute modality architectures are the same for all models. Our models also have very close results to supervised prediction models that are trained to predict the attribute or mask directly from the image modality taken from  Wesego and Rooshenas (2023). Unconditional results also show that the baselines do not generate good-quality images unconditionally. The effect of the auxiliary score-based model can be seen in this case, where MDDVAE can leverage it to generate coherent outputs along all the modalities. In general, the results indicate that our proposed MDDVAE model exhibits superior performance in both conditional and unconditional generation tasks, consistently producing higher-quality and more coherent outputs compared to the baselines.
Finally, in addition to quantitative results, we show qualitative results in Figure  8. The figure shows conditionally generated images for MDDVAE and the baselines. MDDVAE generates higher-quality images that are more coherent with the given conditions.
@@ 4.3 Ablation: Effect of the Auxiliary Score Model
Figure 10: Unconditional generation using MDDVAE when using no auxiliary score model (left) and when using auxiliary score model (right)
In this section, we will look at the importance of using the auxiliary score-based model. By default, multimodal VAEs sample from $\mathcal{N}(0,\mathbf{I})$ to perform unconditional generation. The problem with this is that there could be considerable differences in distribution between the prior and the posterior, which will lead to low-quality samples. Previous unimodal VAE works improved sample quality by training an energy-based model on the latent space Aneja et al. (2021); Pang et al. (2020). The energy-based model will serve as a prior instead of the standard normal distribution and provide samples closer to the posterior, thereby improving generated outputs.  Preechakul et al. (2022) also used a score-based model in the unimodal diffusion autoencoder to generate samples unconditionally. In light of these, we train an auxiliary score-based model to generate prior samples for unconditional multi-modal generation. The auxiliary score-based model will transform the standard normal sample to a sample from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$. Figure  10 shows the effect of using the auxiliary score model. The left side of the figure shows outputs without using the auxiliary score model. Those outputs are inferior compared with the right-side ones, which use the auxiliary score model and generate high-quality outputs.
@@ 5 Conclusion and Discussion
In this paper, we introduce the Multimodal Diffusion Decoder Variational Autoencoder (MDDVAE), which enhances the capabilities of traditional multimodal VAEs by incorporating diffusion decoders, leading to superior performance. Additionally, we implement an auxiliary score-based model to bridge the gap between the prior and posterior distributions, achieving high-quality and coherent results even in unconditional scenarios. A notable limitation of MDDVAE is its higher computational demand during sampling due to the iterative diffusion process, despite the use of DDIM sampling to reduce the sampling time. Nonetheless, our model demonstrates significant performance improvements across multiple challenging datasets, advancing the field of multimodal VAEs and positioning them as a preferred choice for multimodal applications. Future work will focus on optimizing the multimodal model to reduce computational overhead. Our contributions provide a robust framework for future research and development in multimodal VAE generative modeling.

### v2

@@ 3.3 Conditional Sampling
To perform conditional sampling (i.e. generate missing modalities given some modalities) utilizing our model, we can leverage the factorized nature of the posterior distribution. We first sample from the joint posterior distribution conditioned on the available modalities and then use the sampled latent variable to generate samples using the respective generative model for each modality. The posterior will be the product of the observed modalities $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})=\prod_{i\in{o}}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where the subscript $o$ denotes observed modalities. Then, we can generate an unobserved modality ${u}$ from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})$ by feeding it to the appropriate decoder $p(\mathbf{x|z)}$. By following this, we can generate any modality from any other given modality.
@@ 3.4 Unconditional Sampling
Unconditional generation occurs when no specific modality is provided as input. In this scenario, the goal is to produce samples that exhibit coherence across all modalities without relying on any initial conditions. Previous multimodal VAEs perform unconditional generation by first sampling from a $\mathcal{N}(0,\mathbf{I})$ and then feeding the sample to each decoder. Doing that may produce suboptimal results as there is a gap between the posterior and the prior (Aneja et al., 2021). To get better samples, we train an auxiliary score-based diffusion model to sample from the posterior. Figure  2 demonstrates how the score model facilitates unconditional sampling. This auxiliary score model is trained using the diffusion objective but on the latent space $\mathbf{z}$ to generate samples from the posterior. The objective is $\mathbb{E}_{t,\mathbf{z}_{0},\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t)\right\|^{2}$ where $\mathbf{z_{0}}$ is sampled from $q_{\phi}(\mathbf{z}|\mathbf{x})=\prod_{i}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where $\mathbf{x}_{i}$ consists of each modality in $\mathbf{x}$. The main objective of training this auxiliary model is to get high-quality samples from the posterior instead of using $\mathcal{N}(0,\mathbf{I})$ as the starting sample for unconditional generation. To generate samples unconditionally, we first sample $\mathbf{z}$ from the auxiliary score model and use the respective generative model $p_{\theta}(\mathbf{x}_{i}|\mathbf{z})$ of each modality to get the final samples.
@@ 4 Experiments
We perform experiments on two datasets. The first one is from Wah et al. (2011), which contains images of birds with text captions describing the birds. We use actual images of the birds instead of ResNET features, unlike previous multimodal VAEs that struggle with image data (Shi et al., 2019). The second dataset is a high-dimensional CelebAMask-HQ  (Lee et al., 2020) dataset that consists of images, masks, and attributes of celebrities expressed in the three modalities. We compare different multimodal VAE baselines consisting of MVTCAE (Hwang et al., 2021), MoPoE (Sutter et al., 2021), and MMVAE+ (Palumbo et al., 2023). MMVAE+ is not included in the CUB dataset experiment as we use pre-trained VAE for the text captions, which will require changing the structure of the text VAE because of the modality-specific and shared representations in MMVAE+. Nonetheless, we added the results from their paper for the text-to-image results, which can be referenced from their work (Palumbo et al., 2023). We also add additional baselines from  Wesego and Rooshenas (2024); Bounoua et al. (2024) that use a latent score-based diffusion model to learn the joint latent space discussed in the next sections. We compare these baselines with our proposed model variants Diff-MVAE and Diff-MVAE*. Diff-MVAE* is trained using the same objective as eq. 3.2, but with $\lambda(t)=1$. This choice introduces a mismatch between the ELBO and the score-matching objective, resulting in a violation of the ELBO in Eq. 3.2. However, empirical results indicate that setting $\lambda(t)=1$ leads to higher-quality generated samples, a finding that has also been reported in previous works (Ho et al., 2020; Vahdat et al., 2021).
@@ 4.1 Caltech Birds (CUB)
Figure 5: Conditional samples given text using Diff-MVAE (left), Diff-MVAE*(right)
The CUB dataset consists of two modalities: image and text describing the image (caption). In order to get meaningful text outputs, we initialized the text VAE from the pre-trained weight of  Li et al. (2020) that uses a BERT encoder and GPT-2 decoder for all models. The images are resized to 64x64. We used a latent size of $768$ for both the image and text. The encoder architecture for the image modality is similar to that of  Daniel and Tamar (2021) and the diffusion decoder uses a UNET architecture. The models were trained for 500 epochs using the Adam optimizer (Kingma and Ba, 2015). The baseline models are trained using different $\beta$ values listed in the appendix, and the best model was selected by using the average of the conditional and unconditional FID. The unconditional auxiliary score model also uses a 1D UNET architecture and accepts input similar to the size of the latent dimension. DDIM sampling was used for 50 steps to generate samples for the image modality for Diff-MVAE* and for unconditional sampling (Song et al., 2021a). We used Euler–Maruyama sampling for Diff-MVAE. More details about the training and inference are in the Appendix section.
this bird has wings that are brown and has a white belly
a small bird with a long bill and bright orange breast
Figure 7: Unconditional samples using Diff-MVAE (left), Diff-MVAE*(right)
We use FID, which is a widely adopted metric for evaluating the quality of images, as the evaluation metric to measure conditionally and unconditionally generated samples (Heusel et al., 2017). In addition to that, we use Clip-Score to measure the similarity of the image-text pairs  (Hessel et al., 2021; Bounoua et al., 2024). It’s clear that a good multimodal VAE model will have coherent unconditional samples where the image and text represent the same concept expressed in its respective modality. In addition, conditionally generated images should follow the description in the text, and conditionally generated text should follow the given image.
@@
We present the results in Table  1 and  2. In addition to the implemented baselines, we add results from previous works of MLD by  Bounoua et al. (2024) and MMVAE+ on text-to-image FID. The FID scores evaluation demonstrates that our model, Diff-MVAE, outperforms the baselines, generating high-quality images that exhibit greater coherence and alignment with the text modality. In contrast, the baselines struggle to generate meaningful images that are consistent with the textual information, supporting the findings of Daunhawer et al. (2022) regarding the challenges faced by existing multimodal VAEs in difficult tasks. Table  2 also shows that Diff-MVAE generates more coherent conditional outputs that follow the given modality. We also added qualitative results from our model, which is the only multimodal VAE model generating good images in this dataset, are shown in figure  5 and  7 for conditional and unconditional generation respectively.
@@ 4.2 CelebAMask-HQ
The CelebAMask-HQ dataset provides multimodal representations of visual characteristics related to human faces. In this dataset, the images, masks, and attributes can be viewed as distinct yet complementary modalities that capture various aspects of an individual’s appearance. We resize the images and masks to a resolution of 128 by 128 pixels. We combine multiple masks of parts of the face into a single 0/1 mask similar to  Wesego and Rooshenas (2024). Specifically, all the provided masks in the CelebAMask-HQ dataset, except for the skin mask, are drawn on top of each other, resulting in a single composite mask. This composite mask encodes the presence or absence of various facial features. For the attribute modality, we follow the preprocessing approach described by Wu and Goodman (2018); Wesego and Rooshenas (2024). Out of the 40 existing attributes in the CelebAMask-HQ dataset, we selectively choose 18 attributes that are most relevant and informative for representing the visual characteristics of a person’s face. We used a latent size of $256$. Similar to the previous dataset, the encoder and decoder architecture for the image modality are inspired by  Daniel and Tamar (2021) except for the diffusion decoder, which uses a UNET architecture. The unconditional auxiliary score-based model also works on the $256$ dimension latent size resized to 16x16 utilizing a UNET architecture. We used a similar sampling strategy of the previous dataset for Diff-MVAE variants. More details about the experimental setup are located in the appendix section. The baselines here are MoPoE, MVTCAE, MMVAE+ plus SBM-VAE-C  (Wesego and Rooshenas, 2024) and MLD  (Bounoua et al., 2024), which use unimodal variational autoencoders (VAEs) or autoencoders (AEs) for each modality and fuse them using a score-based model for learning the joint distribution. We also add MoPoE*, which is trained with the same hyperparameters as Diff-MVAE but without diffusion decoders.
(a) Diff-MVAE
(b) Diff-MVAE*
(c) MVTCAE
(d) MMVAE+
@@
Diff-MVAE 0.76($\pm$0.001) 0.75($\pm$0.001) 42.3($\pm$0.2) 41.9($\pm$0.06) 45.1($\pm$0.45) 43.5($\pm$0.2) 0.90($\pm$0.001) 0.90($\pm$0.001)
Diff-MVAE* 0.76($\pm$0.001) 0.76($\pm$0.001) 28.3($\pm$0.02) 28.5($\pm$0.1) 32.3($\pm$0.35) 28.4($\pm$0.3) 0.90($\pm$0.001) 0.90($\pm$0.001)
MoPoE* 0.74($\pm$0.002) 0.75($\pm$0.002) 104.8($\pm$0.31) 105.8($\pm$0.18) 182.3($\pm$0.29) 139.8($\pm$0.65) 0.90($\pm$0.001) 0.90($\pm$0.001)
@@
We present our results in table  3 which shows the performance of the models in both conditional and unconditional settings. Diff-MVAE generates high-quality images, achieving the best FID scores among all models in image generation. Not only that but also the performance of the model in the other modalities where Diff-MVAE outperforms the baselines in attribute prediction, even if the attribute modality architectures are the same for all models. Our models also have very close results to supervised prediction models that are trained to predict the attribute or mask directly from the image modality  (Wesego and Rooshenas, 2024). Unconditional results also show that the baselines do not generate good-quality images unconditionally. The effect of the auxiliary score-based model can be seen in this case, where Diff-MVAE can leverage it to generate coherent outputs along all the modalities. In general, the results indicate that our proposed Diff-MVAE model exhibits superior performance in both conditional and unconditional generation tasks, consistently producing higher-quality and more coherent outputs compared to the baselines. We show qualitative results in Figure  9. The figure shows conditionally generated images for Diff-MVAE and the baselines where Diff-MVAE generates higher-quality images that are more coherent with the given conditions.
@@ 4.3 Ablation: Effect of the Auxiliary Score Model
Figure 11: Unconditional generation using Diff-MVAE* when using no auxiliary score model (left) and when using auxiliary score model (right)
This section covers the importance of using the auxiliary score-based model. By default, multimodal VAEs sample from $\mathcal{N}(0,\mathbf{I})$ to perform unconditional generation. The problem with this is there could be considerable differences in distribution between the prior and the posterior, which will lead to low-quality samples. Previous unimodal VAE works improved sample quality by training an energy-based model on the latent space that will serve as a prior instead of the standard normal distribution and provide samples closer to the posterior, thereby improving generated outputs (Aneja et al., 2021).  Preechakul et al. (2022) also used a score-based model in the unimodal diffusion autoencoder to generate samples unconditionally. In light of these, we train an auxiliary score-based model to generate prior samples for unconditional multi-modal generation. The auxiliary score-based model will transform the standard normal sample to a sample from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$. Figure  11 compares outputs with and without the auxiliary score model, showing inferior results on the left and improved ones on the right.
@@ 5 Conclusion and Discussion
In this paper, we introduce Diff-MVAE, a new multimodal VAE that enhances the capabilities of traditional multimodal VAEs by incorporating diffusion decoders, leading to superior performance. Additionally, we implement an auxiliary model to bridge the gap between the prior and posterior distributions, achieving high-quality and coherent results. A notable limitation of Diff-MVAE is its higher computational demand during sampling due to the iterative diffusion process. Nonetheless, our model demonstrates significant performance improvements across multiple challenging datasets, advancing the field of multimodal VAEs and positioning them as a preferred choice for multimodal applications. Future work will focus on optimizing the multimodal model to reduce computational overhead. Our contributions provide a robust framework for future research and development in multimodal VAE generative modeling.